### v2

# The Bose-Hubbard polaron from weak to strong coupling

Abstract
We investigate the zero-temperature properties of a mobile impurity immersed in a bath of bosonic particles confined to a square lattice. We analyze the regimes of attractive and repulsive coupling between the impurity and the bath particles for different strengths of boson-boson interactions in the bath, using exact large-scale quantum Monte-Carlo simulations in the grand canonical ensemble. For weak coupling, the polaron mass ratio is found to decrease around the Mott insulator (MI) to superfluid (SF) transition of the bath, as predicted by recent theory [1], confirming the possible use of the impurity as a probe for the transition. For strong coupling in the MI regime, instead, the impurity is found to modify the bath density by binding to an extra bath particle or a hole, depending on the sign of the polaron-bath interactions. While the binding prevent the aforementioned use of the polaron mass ratio as an MI-SF transition probe, we show that it can be used instead as a probe of the binding itself. Our exact numerical results provide a benchmark for comparing lattice Bose polaron theories and are relevant for experiments with cold atoms trapped in optical lattices, where the presence of a confining harmonic potential can be modeled by a slowly varying local chemical potential.
A mobile impurity interacting with a bath can give rise to a quasiparticle known as a polaron [2, 3], a paradigmatic open quantum system relevant to semiconductors [4], superfluid helium [5], and nuclear matter [6]. Recent advances in ultracold atom experiments, including quantum gas microscopy [7, 8, 9, 10], have enabled detailed studies of impurities in Bose-Einstein condensates [11], both in the continuum and in optical lattices described by the Bose-Hubbard (BH) model [12, 13, 14, 15, 9, 16, 17, 18, 19, 20, 21]. This has opened up the study of polaron physics in the strongly interacting regime and in the vicinity of quantum phase transitions. Lattice Bose polarons have been recently explored near the Mott insulator–superfluid (MI–SF) transition using beyond-mean-field Quantum-Gutzwiller (QGW) methods [1], and in combination with diagrammatic and variational quantum Monte Carlo techniques [22] at fixed particle density and in the canonical ensemble. These works demonstrate that for sufficiently weak impurity-bath interactions, polaron spectral properties such as the energy shift and effective mass can serve as sensitive probes of the MI–SF transition and its universality class [1]. At strong coupling, novel bound states involving the impurity and bath particles-holes excitations have been shown to emerge for fixed atoms numbers, beyond the polaron picture [22]. A key challenge remains in providing exact results for polaron dynamics across all regimes of interaction, in particular for strong impurity-bath and intra-bath interactions and in the grand canonical ensemble. The latter is particularly interesting for neutral atom experiments, where the presence of external harmonic confinement for the atoms effectively realizes a situation where local atomic densities vary across the lattice, providing effective particle reservoirs.
In this work, we use exact large scale quantum Monte Carlo (QMC) simulations based on the worm algorithm [23] to study a mobile impurity immersed in a Bose-Hubbard bath in two dimensions. Our multi-species worm algorithm (see Supplemental Material) can describe all regimes of impurity-bath interactions — from weak to strong, attractive to repulsive — including possible impurity-atom and impurity-hole binding. In order to characterize different regimes of impurity-bath interactions and possible new binding mechanisms, we focus on spectral properties such as the polaron effective mass and the bath-induced shift as well as impurity-bath correlation functions. For weak impurity-bath interactions, the polaron mass ratio is found to decrease around the MI-SF transition and the polaron can be used as a probe of the transition, as predicted by [1]. We quantify for what strength of interactions this picture ceases to describe the system dynamics and demonstrate that for strong interactions the impurity locally modifies the bath density resulting in a binding to an extra bath particle or a hole, depending on the sign and strength of the impurity-bath interactions (see Fig. 1(a)). These binding mechanisms, which tend to suppress the polaron mass ratio and differ from the case of constant density [22], produce distinct spectroscopic signatures. We discuss the connection of these results in the grand-canonical ensemble with neutral atom experiments in the presence of an external harmonic potential.
We consider a system comprising a single mobile impurity coupled to a BH bath on a uniform square lattice with $N=L\times L$ sites, and lattice spacing $a=1$. The microscopic Hamiltonian $H=H_{B}+H_{I}+H_{IB}+H_{\Omega}$ reads
[…]
Here, $H_{B}$, $H_{I}$, $H_{IB}$ and $H_{\Omega}$ represent the Hamiltonians for the bath, the impurity, the bath-impurity interaction, and the external harmonic trap acting on both species, respectively. The chemical potential of the bath is included in $H_{\Omega}$. The bosonic operators ${a}_{i}^{\dagger}$ (${a}_{i}$) create (annihilate) a bath particle at lattice site $i$, with ${n}_{B,i}={a}_{i}^{\dagger}{a}_{i}$ the local density operator. Similarly, the impurity is described by the operators ${b}_{i}^{\dagger}\;({b}_{i})$, which create (annihilate) impurity particles, with ${n}_{I,i}={b}_{i}^{\dagger}{b}_{i}$. The on-site interaction energy and chemical potential of the bath are denoted by $U$ and $\mu$, respectively, while $U_{IB}$ characterizes the interaction strength between the impurity and the bath particles. The parameter $t$ ($\tilde{t}$) represents the nearest-neighbor hopping amplitude for a bath particle (impurity), $r_{i}$ is the distance of site $i$ from the center of the trap, and $\Omega$ represents the trap energy. We focus on impurity-bath interaction effects, assuming equal hopping, $t=\tilde{t}$. We explore scenarios involving both weak and strong impurity-bath couplings, $U_{IB}/t$, for both attractive ($U_{IB}<0$) and repulsive ($U_{IB}>0$) interactions.
Our analysis focuses on two complementary settings: (i) a homogeneous system $(\Omega=0)$ in the grand canonical ensemble, which provides a controlled environment for estimating polaron observables and interaction-driven crossovers; and (ii) a harmonically trapped system $(\Omega\neq 0)$ at fixed total particle number, relevant to experiments, where the local chemical potential varies spatially as $\mu(r)=\mu-\Omega r^{2}$. For sufficiently strong interactions $U/t$, the trapped system exhibits the characteristic “wedding-cake” density profile [24, 25, 26, 27, 28] featuring incompressible MI cores surrounded by compressible SF shells, as shown in Fig. 1(b) and (c). The two settings are connected by the locally grand-canonical nature (ii), where spatial variations in the trap correspond to scanning the chemical potential in the homogeneous case.
Figure 1: (a) Schematic representation of the different states of the system in the MI regime. For $\mu-U<U_{IB}<\mu$, the impurity weakly perturbs the bath and remains mobile. For $U_{IB}<\mu-U$ ($U_{IB}>\mu$), the impurity binds to an extra bath particle (hole), forming a mobile bound pair that propagates via correlated (anti-correlated) hops, respectively (colored arrows). (b) QMC snapshots of a single impurity (red) in a bath of $3$ bosons (blue) in a harmonic trap with $\Omega/t\approx 1.777$ at $2dt/U=0.05$, for $U_{IB}/U=-1$ (left), 0.2 (center), and 0.5 (right). Marker areas are proportional to the local occupations $\langle n_{I}\rangle$ and $\langle n_{B}\rangle$. $\Delta E$ is the energy cost of moving a bath particle from the edge to the trap center.(c) Same parameters as in (b), but marker areas represent the deviation from the impurity-free bath, $\langle n_{B}\rangle-\langle n_{B,\text{pure}}\rangle$; green (purple) indicates enhanced (reduced) bath density.
We begin by analyzing the polaron at $\mu=(\sqrt{2}-1)U$, which corresponds to the mean-field tip of the first Mott lobe and allows for direct comparison with recent theoretical studies [1, 22]. For reference, the MI–SF transition in the 2D Bose–Hubbard model at unit filling along the commensurate density line has been determined via QMC to occur at $2dt/U=0.2389(6)$ [29]. In contrast to variational QMC results obtained in the canonical ensemble [22], our numerically exact simulations in the grand-canonical ensemble (see Supplemental Material) show that a single impurity does not shift the transition point or affect the Mott phase properties—such as the excitation gap or compressibility—in the thermodynamic limit.
Figure 2: Top: Polaron energy shift $E_{0}$ as a function of impurity-bath interaction strength $U_{IB}/U$ across different bath phases: (a) Mott Insulator (MI) $(2dt/U=0.05)$, (b) near the MI-SF transition (T) $(2dt/U=0.225)$, and (c) superfluid (SF) $(2dt/U=0.35)$. Solid points represent QMC estimates of $E_{0}$, while dashed lines show theoretical shifts. $E_{+1}=2U_{IB}+U-\mu$. Insets: Impurity-bath density-density correlation $C_{IB}(r=0)$, scaled by $L^{2}$, showing local bath occupation changes. Bottom: effective mass ratio $M/M^{*}$ as a function of $U_{IB}/U$, illustrating how impurity-bath interactions modify the impurity’s effective mass in different bath phases: (d) Mott Insulator (MI), (e) near the MI-SF transition (T), and (f) superfluid (SF). All plots are for an $8\times 8$ lattice. Vertical dashed lines indicate $U_{IB}/U$ values at which energy level crossings occur. Error bars that are not visible are within the marker size.
We characterize the quasiparticle properties of the polaron by analyzing its energy dispersion at small momenta, $E_{p}(\boldsymbol{k})=E_{0}+\frac{\boldsymbol{k}^{2}}{2M^{*}}+\mathcal{O}(\boldsymbol{k}^{4})$. Here, $E_{0}=E_{p}(\boldsymbol{k}=0)$ denotes the interaction-induced energy shift, while $M^{*}$ is the polaron’s effective mass, reflecting its mobility within the bath. To estimate $M^{*}$, we compute the impurity’s diffusion coefficient $D$ from its mean-squared displacement between imaginary times $\tau=0$ and $\tau=\beta/2$, and use the the relation $M^{*}=1/(2D)$ [30, 31]. Specifically,
$D=\lim_{\beta\to\infty}\left\langle\left(r(0)-r\left(\frac{\beta}{2}\right)\right)^{2}\right\rangle/\left(2d\cdot\frac{\beta}{2}\right),$ (2)
where $r(\tau)$ denotes the impurity’s position at imaginary times $\tau$, and $d=2$ is the dimensionality of the system. The effective mass $M^{*}$ is then used to estimate the bath-induced shift $E_{0}$ via the relation
$E_{0}=E_{\text{B+I}}-E_{B}-E_{M^{*}},$ (3)
where $E_{\text{B+I}}$ ($E_{B}$) is the total energy of the system in the presence (absence) of the impurity, and $E_{M^{*}}=-d/M^{*}$ is the ground state energy of a free lattice boson of mass $M^{*}$. We note that Eq. (3) includes $O(\boldsymbol{k}^{4})$ contributions in $E_{0}$, negligible in the quasi-particle picture. Moreover, the grand canonical ensemble allows the presence of the impurity to modify the bath’s particle count with respect to a pure bath. Such a modification would be reflected by a modification of the chemical potential contribution in the total energy, and thus would also be taken into account in the value of $E_{0}$. We choose to use this definition as the chemical potential contribution would also appear in the case of a trapped system, even with a fixed number of particles, as discussed later. We also compute the same-site bath-impurity density-density correlation $C_{IB}(r=0)=\left\langle n_{B,i}n_{I,i}\right\rangle_{i}-\left\langle n_{B,i}\right\rangle_{i}\left\langle n_{I,i}\right\rangle_{i}$, with $\left\langle\right\rangle_{i}$ a site and thermodynamic average. This observable compares the number of bath bosons at the impurity site with the site-average number of bath bosons in the system.
Impurity spectral properties and bound states— Figure 2 shows QMC data for the bath-induced energy shift $E_{0}$ (panels (a-c)), the onsite impurity-bath correlation $C_{IB}(0)$ (insets), and the polaron mass ratio $M/M^{*}$ (panels (d-f)), with $M=1/(2t)$. Results are shown as a function of $U_{IB}/U$ (both attractive and repulsive) for three bath regimes: deep in the MI phase (panels (a), (d)), within the MI phase but close to the MI-SF transition (panels (b), (e)), and in the SF phase (panels (c), (f)). In particular, panel (a) shows that, for a bath in the deep MI phase, the bath-induced shift $E_{0}$ takes the values $E_{0}\simeq E_{+1}=2U_{IB}+U-\mu$ (dashed dotted line), $U_{IB}$ (dashed line) and $\mu$ (dotted line) for increasing $U_{IB}/U$ in the regions $U_{IB}<\mu-U$, $\mu-U<U_{IB}<\mu$, and $U_{IB}>\mu$, respectively. These regions correspond to the configurations shown in Fig. 1(a): an impurity bound to an extra boson ($\text{MI}_{+1}$, left), a free polaron (MI, center), and an impurity bound to a hole ($\text{MI}_{-1}$, right). In particular, for small values of $U_{IB}/U$ in the MI phase, $E_{0}\simeq U_{IB}$, consistent with a mean-field picture of the impurity interacting with a homogeneous bath at $\langle n_{B}\rangle=1$. For strong repulsion, $E_{0}\simeq\mu$ reflects the energy cost of removing a bath particle, while for strong attraction, $E_{0}\simeq 2U_{IB}+U-\mu$ accounts for the addition of a boson and its interaction with the impurity and the existing bath particle. The picture above is corroborated by $C_{IB}(0)$, which varies from +1 to 0 to -1 across $\text{MI}_{+1}$, MI and $\text{MI}_{-1}$ regions, respectively. Notably, the extra hole or particle on the impurity site reflects a true change in the bath particle number, rather than a particle–hole pair excitation (see Supplemental Material). This contrasts with Ref. [22], where the impurity cannot alter the bath occupation.
Figure 3: Polaron mass ratio $M/M^{*}$ as a function of $2dt/U$ for (a) attractive and (b) repulsive branches. Different curves correspond to varying impurity-bath interaction strengths $U_{IB}/U$. Insets: zooms on black dotted boxes, showing results consistent with [1]. Dashed grey vertical line: MI-SF transition point of the pure bath (no impurity present). Dashed black line: guide for the eye of the transition between the polaron and dimer regimes for strong repulsive interaction. All plots are for a $8\times 8$ lattice.
Panel (b) shows qualitatively similar features to (a), except for small deviations from the values $E_{+1}$, $U_{IB}$ and $\mu$ near the transition points between the different regions $\text{MI}_{+1}$, MI, and $\text{MI}_{-1}$. The transitions are instead completely smoothed out for the case of a weakly interacting bath in the SF regime (panel (c)), where $E_{0}/U$ increases smoothly and monotonically from attractive to repulsive $U_{IB}/U$. The values of $E_{0}/U$ are qualitatively accounted for by a continuous change in bath density on the site of the impurity and in the polaron cloud in this regime (see Supplemental Material).
The three regions are also clearly identified from the mass ratio $M/M^{*}$ in panel (d-f). In particular, for a bath prepared well into the MI phase (panel (d)) and for $0.5\lesssim U_{IB}/U\lesssim 0.5$, $M/M^{*}$ is close to unity, consistent with the results of [1] for small $U_{IB}$. In the $\text{MI}_{\pm 1}$ regions, instead $M/M^{*}\simeq 0$, which is consistent with our picture above of an impurity bound to a particle and a hole, respectively. Indeed, on a 2D lattice, a free dimer consisting of two similar bosons bound by $U_{IB}$ is expected to have an effective hopping energy $2t^{2}/|U_{IB}|$ for $|U_{IB}|\gg t$ [32, 33], resulting in an expected $M/M^{*}\simeq 10^{-2}$, which is consistent with our QMC results. For panel (e), the transitions are again smoothed out, and the mass ratio is higher in the $\text{MI}_{\pm 1}$ regions, as expected from the higher value of $2dt/U$. Panel (f) shows a continuous evolution of $M/M^{*}$ from strong attractive to strong repulsive interactions for a weakly interacting bath in the SF phase. The characterization of impurity spectral properties across all interaction regimes and in particular the demonstration of dimer states in the $\text{MI}_{\pm 1}$ regimes in the grand canonical ensemble are main results of this work.
MI-SF transition— Figure 3 shows the polaron mass ratio $M/M^{*}$ versus $2dt/U$ for various values of $U_{IB}/U$. In the weak coupling limit — darker circles and squares in the attractive (panel (a)) and repulsive (panel (b)) branches — the ratio stays close to unity, consistent with the picture of a mean-field-type free polaron deep inside the MI phase. As $2dt/U$ approaches the MI-SF boundary (vertical grey dashed line), $M/M^{*}$ exhibits a pronounced dip (see Inset), whose depth increases with $|U_{IB}/U|$. Beyond the transition, the mass ratio rises again in the SF phase. This non-monotonic “dip-rebound” behavior is consistent with quantum Gutzwiller (QGW) predictions [1], and serves as a clear spectral signature of the MI-SF transition in the weak coupling regime.
As $|U_{IB}/U|$ increases beyond this regime, the polaron evolves from a mobile quasiparticle into a bound dimer in the MI phase. The mass ratio $M/M^{*}$ then displays two distinct behaviors depending on $2dt/U$. At low hopping, the impurity binds strongly to a particle or hole excitation in the bath, forming a heavy dimer-like state. This is reflected in the strongly suppressed mass ratio observed for $|U_{IB}/U|\gtrsim 0.6$ at low $2dt/U$, signaling the $\text{MI}_{\pm 1}$ region. As $2dt/U$ increases, the bound state destabilizes, and the system crosses over into the conventional MI phase, where the impurity regains mobility. This crossover, marked by curved dashed black line in the figure, occurs at progressively larger $2dt/U$ as the $|U_{IB}/U|$ increases, indicating a gradual strengthening of the bound state. This interpretation is further supported by the accompanying changes in bath density and same-site correlation $C_{IB}(0)$ (see Supplemental Material). These results show that the mass ratio ceases to serve as a reliable probe of the MI–SF transition in the strong-coupling regime. However, it remains a robust indicator of impurity–bath bound-state formation.
Figure 4: (a) Polaron energy $E_{p}/U$, (b) bath-induced shift $E_{0}$ normalised by the mean field shift $U_{IB}\langle n_{b}\rangle$, (c) mass ratio $M/M^{*}$ vs $\mu/U$ for $U_{IB}/U=0.1$ (solid blue line, circular markers) and 0.1-0.1 (dashed red line, square markers). Dashed vertical lines represent the MI-SF transition points of the bath based on estimated bath density and superfluid fraction. Dotted vertical lines are guides for the eyes for the position of the sharp features. Here, $2dt/U=0.05$. All plots are for an $8\times 8$ lattice. The $\mu/U$ scale is a pseudo-log10 centered at 0.5 and scaled by a factor 0.3.
Harmonic trap and experimental observables— Having established that increasing $|U_{IB}/U|$ in a homogeneous bath drives the polaron from a mobile quasiparticle to a bound dimer, we now ask how this physics manifests in a harmonically confined system. In the presence of a trap, spatially varying chemical potential, together with strong on-site interactions drive the bath into the familiar “wedding-cake” density profile [24, 26, 27, 28]. Figure 1(b) shows QMC snapshots of bath and impurity densities for varying impurity–bath couplings $U_{IB}$. When the bath is initially prepared in an MI state with a thin SF shell, a sufficiently strong $U_{IB}$ locally distorts the bath density—enhancing or depleting it depending on the sign of $U_{IB}$—with compensating changes occurring at the bath boundaries (see Fig.1(c)). Although the total number of bath particles is fixed, the compressible SF shell acts as an effective reservoir, allowing for local particle-number fluctuations. In this sense the system is locally grand-canonical: moving a bath particle from radius $r_{1}$ to $r_{2}$ costs $\Delta E=-\Omega(r_{1}^{2}-r_{2}^{2})$, which plays the role of a local chemical-potential shift. For example, in Figure 1(b), $\Delta E/U\approx-0.2$ for moving a particle from the outer edge to the center.
Since the harmonic confinement effectively lowers the local chemical potential with increasing distance from the trap center, we simulate this radial variation by fixing $2dt/U$ and scanning $\mu/U$. Figure 4 shows QMC results for three key observables: the mass ratio $M/M^{*}$ (panel a), the bath-induced shift $E_{0}$ normalized by the mean-field value $U_{IB}\langle n_{B}\rangle$ (panel b), and the polaron energy $E_{p}=E_{B+I}-E_{B}$ (panel c), all as functions of $\mu/U$. These are shown for weak impurity–bath interactions $U_{IB}/U=\pm 0.08$ at fixed $2dt/U=0.05$. In the range $0.08\lesssim\mu/U\lesssim 0.88$, all observables remain nearly constant, corresponding to the regime where the bath is in a uniform Mott insulating state with $\langle n_{B}\rangle=1$. Outside this interval, sharp features emerge. For $\mu\simeq 0.08$, the repulsive polaron (blue solid line, circular markers) exhibits a sharp cusp in $M/M^{*}$ and a pronounced drop in $E_{0}$ near $\mu/U\simeq U_{IB}/U$ in the MI phase, while the attractive polaron (red dashed line, square markers) shows smoother variations near the MI–SF boundary of the bath (indicated by vertical dashed line). Conversely, at large $\mu\simeq 0.88$, it is the attractive polaron that exhibits a sharp increase in $E_{0}$ and a corresponding cusp in $M/M^{*}$ near $\mu/U\simeq 1-|U_{IB}|/U$, whereas the repulsive case remains smooth across the MI-SF boundary. These sharp transitions, which are not aligned with the MI–SF boundaries, are consistent with a crossover from the MI phase to the impurity-bound $\text{MI}_{\pm 1}$ phases discussed earlier. The sudden suppression of $M/M^{*}$ signals the formation of a tightly bound bath–impurity dimer. Interestingly, the polaron energy $E_{p}$ (panel c) does not show corresponding discontinuities—its behavior remains smooth—implying that the abrupt changes in $M/M^{*}$ and $E_{0}$ cancel each other out in $E_{p}$.
In a harmonic trap, low values of the chemical potential $(\mu\lesssim 0.08)$ correspond to the outer edge of the trap and represents the superfluid (SF) shell, where the bath density $\langle n_{B}\rangle<1$. In contrast, high chemical potentials $(\mu\gtrsim 0.9)$ correspond to the SF region between the $\langle n_{B}\rangle=1$ and $\langle n_{B}\rangle=2$ Mott insulating (MI) plateaus, as seen in the typical “wedding-cake” structure [26, 27, 8]. These results indicate that impurity binding can occur even at weak impurity–bath coupling in a trapped system, and not just in the strong-coupling regime previously discussed in Fig. 1(b) and (c). The onset of binding is clearly reflected in the sharp changes observed in effective mass $M^{*}$ and bath-induced energy shift $E_{0}$, providing experimentally accessible signatures. In practice, by tuning the harmonic confinement—trap frequencies/curvature, depth, aspect ratio, and center offset—as well as the total atom number, one can engineer the desired bath phase at the trap center [28]; meanwhile, the impurity can be positioned independently with species-selective optical tweezers [34, 35], enabling precise placement within targeted regions of the inhomogeneous bath.
In conclusion, we demonstrated through exact numerical analysis that increasing the impurity-bath coupling strength in a lattice Bose polaron—defined by a single impurity boson interacting with a Bose-Hubbard bath in a grand canonical ensemble to the emergence of an MI-like phase with incommensurate density. This occurs for both attractive and repulsive interactions, as the impurity binds an additional bath particle or hole. Moreover, we demonstrated that the properties of the polaron, such as its effective mass, exhibit distinct behaviors near the MI-SF phase transition, depending on whether the impurity-bath coupling is weak or strong. We showed that the polaron mass ratio can either be used as a probe for the bath’s MI-SF phase transition for weak coupling, or as a probe of the $\text{MI}_{\pm 1}$ bound state for strong coupling. We argued that these results are relevant for the experimental study of the Bose lattice polaron in an harmonic trap, where the added spatial complexity make those effects possible even for a fixed number of bosons and weak or strong impurity-bath interaction. This work opens several research directions, including the study of bipolaron formation and many-polaron problems in the strong interacting regime.
Acknowledgements.
Acknowledgments: We gratefully acknowledge discussions with Ragheed Alhyder and Chao Zhang. This research has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie project 955479 (MOQS), the Horizon Europe programme HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa) and from the French National Research Agency under the Investments of the Future Program projects ANR-21-ESRE-0032 (aQCess), ANR-22-CE47-0013-02 (CLIMAQS) and ANR-23-CE30-50022-02 (SIX).